# Projection-free computation of robust controllable sets with constrained zonotopes

### v1

@@ 7 Application: Abort-safe rendezvous
Abort safety in spacecraft rendezvous [13, 29, 11] requires that a spacecraft in nominal operation approaches a target while retaining the ability to avoid collision with the target in the event of an anomaly or a failure. In [11], we showed that the problem of abort-safe spacecraft rendezvous could be encoded using RC sets, and we computed these sets using polytopes. However, such an approach is unreliable in high-dimensions and suffers from numerical issues, which motivates the computation of the RC sets using constrained zonotopes. Additionally, in order to guarantee safety, we require an exact computation or inner-approximation of the RC sets.
In this work, we consider rendezvous to a future Lunar gateway flying in a near-rectilinear halo orbit (NRHO) around the Moon [18]. To minimize the use of fuel, we allow the spacecraft to utilize all 3 degrees of freedom in its approach. We compute six-dimensional RC sets that constrain the rendezvous trajectory in order to guarantee that, in the event of an anomaly, off-nominal operation of the spacecraft allows for a safe abort maneuver. Additionally, the trajectory must lie in a line-of-sight cone that arises from sensing and communication requirements, and is determined to ensure that the Sun stays behind the spacecraft, maximizing the perceptibility of the Lunar gateway.
@@
with state (position and velocity) $x_{t}\in\mathbb{R}^{6}$, input $u_{t}\in\mathcal{U}\subset\mathbb{R}^{3}$, $B_{t}$ as the last three columns of $A_{t}$, and perfect state measurements. In (58), the input $u_{t}$ corresponds to impulsive changes in velocities.
Off-nominal dynamics: We consider three modifications to the nominal dynamics in the event of failure — limited actuation to model the event where the main thrusters fail and the spacecraft is forced to use redundant thrusters like attitude thrusters, process noise to model the resulting actuation uncertainty, and measurement noise to model sensing uncertainty that may increase with the use of redundant thrusters. The need for redundant thrusters is well-known in space applications to ensure safety in off-nominal scenarios [31]. We assume that the process and measurement noises are drawn from pre-determined bounded sets that may be characterized via offline statistical analysis [31].
@@
where $z_{k|t}$ is the state after failure at time $k\geq t$, the failure trajectory is initialized by $z_{t|t}=x_{t}$ the nominal state at failure time $t$, and $w_{k|t}\in\mathcal{W}_{\text{off-nom}}\subset\mathbb{R}^{3}$ and $\eta_{k|t}\in\mathcal{E}_{\text{off-nom}}\subset\mathbb{R}^{6}$ are bounded disturbances to the input and post-failure state respectively. The bounded disturbances model the actuation mismatch and sensing limitations that can become prominent after failure. Additionally, the post-failure input $u_{t}\in\mathcal{U}_{\text{off-nom}}\subset\mathcal{U}$, where $\mathcal{U}_{\text{off-nom}}$ encodes the limited actuation available after failure. Since we have access to only imperfect measurements post-failure, we consider a feedback controller $\pi:\mathbb{R}^{6}\to\mathcal{U}_{\text{off-nom}}$ that provides a post-failure control $u$ in (7) given the current state estimate $\hat{z}\in\mathbb{R}^{6}$. Let $\Pi$ be the set of all such feedback controllers.
@@
with disturbance $\phi_{t}\in\Phi_{t}=\mathcal{E}_{\text{off-nom}}\oplus(B_{t}\mathcal{W}_{\text{off-nom}})\oplus(-A_{t}\mathcal{E}_{\text{off-nom}})$. From (59b), the post-failure true state $z_{t}$ and corresponding state estimate $\hat{z}_{t}$ satisfy
@@
Rendezvous constraints: We consider the problem of navigating the spacecraft to a target set $\mathcal{T}\subset\mathbb{R}^{3}$ in front of the Lunar gateway. The designed nominal rendezvous must lie inside a line-of-sight cone $\mathcal{L}\subset\mathbb{R}^{3}$ originating from the Lunar gateway, and the trajectory must stay outside a keep-out set $\mathcal{D}\subset\mathbb{R}^{3}$ around the Lunar gateway during the rendezvous maneuver. Additionally, for some pre-determined post-failure safety horizon $T_{\text{safe}}\in\mathbb{N}$, at any time $t$, the nominal state $x_{t}$ must satisfy the abort-safety requirement:
@@
For $\lambda>0$, the objective of (69) balances the typical goals of rendezvous — approaching the target set $\mathcal{T}$ while limiting the energy spent. We measure the energy spent in rendezvous [13] as ${(\Delta v)}_{t}={\|u_{t}\|}_{2}$.
Enforcement of (Abort-safety) using RC sets: Similarly to [11], we encode the abort-safety requirements using appropriately defined RC sets. Using (60) and (61), we derive a sufficient condition for (64) based on RC sets. Let $\mathcal{S}^{\complement}=\mathbb{R}^{3}\setminus\mathcal{S}$ be the complement of a set $\mathcal{S}\subseteq\mathbb{R}^{3}$.
@@
See Appendix A.5 for the proof. Motivated by Proposition 6, we solve the optimal control problem (69) using a receding horizon framework for tractability.
For a finite planning horizon $T_{\text{plan}}\in\mathbb{N}$, consider the following (non-convex) optimization problem that approximates (69),
[…]
@@
where $\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})=\mathcal{K}(k,k+T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})\ominus\mathcal{E}_{\text{off-nom}}$. The optimization problem (76) is non-convex due to disjunctive constraints [32].
@@
For a sampling time $T_{\text{sample}}=20$ minutes, we solve (69) using a receding horizon framework with a planning horizon of $2$ hours ($T_{\text{plan}}=6$) and an abort-safety time horizon of $6$ hours ($T_{\text{safe}}=18$). We consider an impulse-based control with the nominal control set $\mathcal{U}=\llbracket-1/3,1/3\rrbracket$ (in m/s), off-nominal input set $\mathcal{U}_{\text{off-nom}}=0.1\mathcal{U}$, post-failure process noise in the ellipsoid $\mathcal{W}_{\text{off-nom}}=(1/60I_{3},0_{3\times 1})$ (in m/s), and post-failure measurement noise in the ellipsoid $\mathcal{E}_{\text{off-nom}}=(\mathrm{diag}(1,1,1,1/60,1/60,1/60),0_{6\times 1})$ (in m and m/s). We define the origin-centered keep-out set $\mathcal{D}={\llbracket-100,100\rrbracket}^{3}$ (in m) which contains the Lunar gateway [18]. We also define a proper cone characterized by four rays originating from the origin as the “line-of-sight cone” =$\mathcal{L}=\{x\in\mathbb{R}^{3}|[0,1,-1;0,1,1;-1,1,0;1,1,0]x\leq 0_{4\times 1}\}$ for ease of discussion. We define an ellipsoidal target set $\mathcal{T}=(0.05I_{3},c_{\text{target}})$ with $c_{\text{target}}=[0.2416;-0.4017;-0.1738]$ and initial state $x_{0}=[1.4498;-2.4105;-1.0429;0.01;0.01;0.01]$ such that the target is $0.5$ km and the initial state is $3$ km away from the origin with non-zero initial velocity. We rotate $\mathcal{D}$ and $\mathcal{L}$ to have the $+\mathrm{y}$-face of $\mathcal{D}$ and the axis of symmetry of $\mathcal{L}$ be aligned with the line segment joining $x_{0}$ and $c_{\text{target}}$ respectively. The choice of parameters considers a rendezvous approach with the Sun behind the spacecraft, as the gateway flies near the apolune of the NRHO.
@@
The exact computation of $\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})$ based on polytopes [4] is unreliable, due to the complexity of the calculations involved in the considered problem setting. Therefore, we use the proposed approach for inner-approximating $T_{\text{safe}}$-RC set using constrained zonotopes to enforce abort-safety constraint. Specifically, we use Theorem 1 and Section 5.2 to compute constrained zonotopic, inner-approximations of $\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})$, and then use big-M formulations to cast the disjunctive constraints in (76) as mixed-integer linear constraints.
@@
Consider $N_{D}$ constrained zonotopes $\{\mathcal{C}_{i}\}_{i=1}^{N_{D}}$ where $\mathcal{C}_{i}\subseteq\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})$ for each $i\in\mathbb{N}_{[1:N_{D}]}$ and some $k\in\mathbb{N}_{[t:t+T_{\text{plan}}]}$. Using $N_{D}$ auxiliary continuous variables $\xi_{i}\in\mathbb{R}^{N_{C,i}}$ and $N_{D}$ auxiliary binary variables $\delta_{i}\in\{0,1\}$, and for any sufficiently large $M>0$, the following set of $12(n+N_{C,i})+M_{C,i}+$ 1 mixed-integer linear constraints is equivalent to the constraint $x_{k|t}\in\cup_{i=1}^{N_{D}}\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})$ at any $k\in\mathbb{N}_{[t+1:t+T_{\text{plan}}]}$,
$\forall i\in\mathbb{N}_{[1:N_{D}]},$ $\quad{\|G_{C,i}\xi_{i}+c_{C,i}-x_{k|t}\|}_{\infty}\leq M(1-\delta_{i})$
$\forall i\in\mathbb{N}_{[1:N_{D}]},$ $\quad A_{C,i}\xi_{i}=b_{C,i},\ {\|\xi_{i}\|}_{\infty}\leq 1,$
$\quad\sum\nolimits_{i=1}^{N_{D}}\delta_{i}\geq 1.$
@@
See [32] for more details on mixed-integer reformulations of disjunctive constraints. Similar mixed-integer constraints based on big-M can be used to encode the disjunctive constraint $x_{k|t}\in\cup_{i=1}^{N_{D}}\mathcal{H}_{i}^{\complement}$ [32]. Thus, we can cast (76) as a mixed-integer quadratic program, and solve it by off-the-shelf solvers like GUROBI [28].
@@
Figure 5: Designed nominal rendezvous trajectory along with the abort trajectories in case of failures at $t\in\mathbb{N}_{[1:T_{\text{sim}}]}$. The nominal trajectory starts at the initial state (diamond) and reaches the target set $\mathcal{T}$ (green) in $T_{\text{sim}}=10$ time steps, while staying within the line-of-sight cone $\mathcal{L}$ (black). The abort-safety requirement curves the nominal trajectory away from the keep-out set (red) at all times. The abort trajectories stay outside the keep-out set, despite the presence of disturbances which are adversarially chosen according to (7).
@@
Using the sets defined in Section 5, we can also compute an abort-safe control $u_{k|t}=\pi_{k}(\hat{z}_{k|t})$ for the current state estimate $\hat{z}_{k|t}$ by solving the convex optimization problem,
$\underset{u_{k|t}}{\text{minimize}}$ $\quad J(u_{k|t})$ (77a)
subject to $\quad u_{k|t}\in\mathcal{U}_{\text{off-nom}},$ (77b)
$\quad A_{t}\hat{z}_{k|t}+B_{k|t}u_{k|t}\in\mathcal{K}_{k+1-t},$ (77c)
@@
where $J$ is a user-specified, convex cost function on the abort-safe control (we choose $J=\|\cdot\|_{2}$ to minimize post-failure fuel consumption), $\{\mathcal{K}_{s}\}_{s=0}^{T_{\text{safe}}}$ is the sequence of sets obtained using Section 5.2 with $\mathcal{K}_{T_{\text{safe}}}=\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}}$ and $\mathcal{K}(t,T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})=\mathcal{K}_{0}$, as prescribed in Proposition 6. We compute the corresponding adversarial disturbance $\phi_{k|t}$ by solving another convex optimization problem,
$\underset{\hat{z}_{k+1|t},\phi_{k|t}}{\text{min}}$ $\quad\mathrm{dist}(\hat{z}_{k+1|t},\mathcal{D})$ (78a)
s. t. $\quad\phi_{k|t}\in\Phi_{k},$ (78b)
$\quad\hat{z}_{k+1|t}=A_{t}\hat{z}_{k|t}+B_{k|t}\pi_{k}(\hat{z}_{k|t})+\phi_{k|t}.$ (78c)
@@
Here, (7) computes an adversarial disturbance that brings the next state estimate as close as possible to the keep-out set. By (7), we approximate the true worst-case disturbance sequence, whose exact computation would have required solving a two-player game, which is computationally difficult in six dimensions [5].
@@
Figure 5 demonstrates the designed nominal rendezvous trajectory, which takes $T_{\text{sim}}=10$ time steps ($200$ minutes) to reach $\mathcal{T}$ with a total $\Delta v$ of $1.07$ m/s. The trajectory design required the computation of $385$ $18$-step RC sets, and the set computation took a total of $7.53$ seconds. Figure 5 also shows the designed abort-safe trajectories originating from each time step of the nominal rendezvous trajectory using the post-failure control (7) and adversarial disturbances (7), along with an outer-approximation of the one-step forward reach set $A_{t}\hat{z}_{k|t}+B_{k|t}\pi_{k}(\hat{z}_{k|t})+\Phi_{k}$. As expected, the abort-safety requirement (64) is satisfied at all times $t\in\mathbb{N}_{[1:T_{\text{sim}}]}$. As the spacecraft approaches the target, we observe that the conservativeness of our enforcement of the abort-safety requirement (64) via Proposition 6 causes the abort trajectories to cluster in front of the keep out set $\mathcal{D}$. In actual rendezvous missions, the spacecraft would receive a go/no-go decision for docking as it nears the target.
@@
Figure 6 shows the evolution of the distance to the target set (in log-scale) and the cumulative $\Delta v$ expended over the course of the rendezvous, as well as the computation time spent solving (76) at each time step $t\in\mathbb{N}_{[0:T_{\text{sim}}]}$. The rendezvous trajectory uses moderately high control inputs $\Delta v$ initially to steer the spacecraft towards the target while satisfying the abort-safety requirements, and then utilizes the momentum to reach the target set with minimal additional control inputs, as expected. The rendezvous trajectory maintains abort-safety using the constrained zonotope-based constraints computed with the method proposed in this paper.

### v2

@@ 7 Application: Abort-safe rendezvous
Abort safety in spacecraft rendezvous [13, 31, 11] requires that a spacecraft in nominal operation approaching a target must retain the ability to avoid collision with the target in the event of an anomaly or a failure. In [11], we showed that the problem of abort-safe spacecraft rendezvous could be encoded using RC sets, and we computed these sets using H-Rep/V-Rep polytopes. However, such an approach is challenging in high-dimensions and suffers from numerical issues, which motivates the computation of the RC sets using constrained zonotopes. Additionally, to guarantee safety, we require an exact computation or inner-approximation of the RC sets.
In this work, we consider rendezvous to a future Lunar gateway flying in a near-rectilinear halo orbit (NRHO) around the Moon [19]. To minimize the use of fuel, we allow the spacecraft to utilize all 3 degrees of freedom in its approach. We compute six-dimensional RC sets that constrain the rendezvous trajectory in order to guarantee that, in the event of a failure, off-nominal operation of the spacecraft allows for a safe abort maneuver. Additionally, the trajectory must lie in a line-of-sight cone that arises from sensing and communication requirements, and ensures that the Sun stays behind the spacecraft to help in perception of the Lunar gateway.
@@
with state (position and velocity) $x_{t}\in\mathbb{R}^{6}$, input $u_{t}\in\mathcal{U}\subset\mathbb{R}^{3}$ models impulsive changes in velocities, and perfect state measurements.
Off-nominal dynamics: We consider three modifications to the nominal dynamics in the event of failure — 1) limited actuation to model the event where the main thrusters fail and the spacecraft is forced to use redundant thrusters like attitude thrusters, 2) process noise to model the resulting actuation uncertainty, and 3) measurement noise to model sensing uncertainty that may increase with the use of redundant thrusters. The need for redundant thrusters is well-known in space applications to ensure safety in off-nominal scenarios [33]. We assume that the process and measurement noises are drawn from pre-determined bounded sets that may be characterized via offline statistical analysis [33].
@@
where $z_{k|t}$ is the state after failure at time $k\geq t$ initialized by $z_{t|t}=x_{t}$ (the nominal state at failure time $t$), and $w_{k|t}\in\mathcal{W}_{\text{off-nom}}\subset\mathbb{R}^{3}$ and $\eta_{k|t}\in\mathcal{E}_{\text{off-nom}}\subset\mathbb{R}^{6}$ are bounded disturbances to the input and post-failure state respectively. The disturbances model the actuation mismatch and sensing limitations that can become prominent after failure. The post-failure input $u_{t}\in\mathcal{U}_{\text{off-nom}}\subset\mathcal{U}$ where $\mathcal{U}_{\text{off-nom}}$ models the limited actuation available after failure. We consider a feedback controller $\pi:\mathbb{R}^{6}\to\mathcal{U}_{\text{off-nom}}$ that provides a post-failure control $u$ in (65) given the current state estimate $\hat{z}\in\mathbb{R}^{6}$. Let $\Pi$ be the set of all such controllers.
@@
with disturbance $\phi_{t}\in\Phi_{t}=\mathcal{E}_{\text{off-nom}}\oplus(B_{t}\mathcal{W}_{\text{off-nom}})\oplus(-A_{t}\mathcal{E}_{\text{off-nom}})$. From (65b), $z_{t}$ and $\hat{z}_{t}$ satisfy
@@
Rendezvous constraints: We consider the problem of navigating the spacecraft to a target set $\mathcal{T}\subset\mathbb{R}^{3}$ in front of the Lunar gateway, while staying inside a line-of-sight cone $\mathcal{L}\subset\mathbb{R}^{3}$ originating from the Lunar gateway. The designed nominal rendezvous trajectory must also stay outside a keep-out set $\mathcal{D}\subset\mathbb{R}^{3}$ around the Lunar gateway during the rendezvous maneuver. Also, for some pre-determined post-failure safety horizon $T_{\text{safe}}\in\mathbb{N}$, the nominal state $x_{t}$ at any time $t$ must satisfy the abort-safety requirement:
@@
For $\lambda>0$, (77) balances the typical goals of rendezvous — approaching the target set $\mathcal{T}$ while limiting the energy spent. We measure the energy spent as ${(\Delta v)}_{t}={\|u_{t}\|}_{2}$.
Enforcement of (Abort-safety) using RC sets: Similarly to [11], we encode the abort-safety requirements using appropriately defined RC sets. Let $\mathcal{S}^{\complement}=\mathbb{R}^{3}\setminus\mathcal{S}$ be the complement of a set $\mathcal{S}\subseteq\mathbb{R}^{3}$.
@@
(Sufficient condition for Abort-safety) Consider a H-Rep polytope keep-out set $\mathcal{D}=\cap_{i=1}^{{L_{D}}}\mathcal{H}_{i}$ with ${L_{D}}$ halfspaces $\mathcal{H}_{i}\subset\mathbb{R}^{3}$, and $\mathcal{E}_{\text{off-nom}}$ that is symmetric about the origin. For any time $s\in\mathbb{N}$, let $\mathcal{K}(s,T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})$ denote the $T_{\text{safe}}$-step RC set for dynamics (66) characterized by ${\{(A_{t},B_{t},\Phi_{t})\}}_{t=s}^{s+T_{\text{safe}}}$, $\mathcal{U}_{\text{off-nom}}$, and $\mathcal{X}_{t}=\mathcal{G}=\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}}$. Then, $x_{s}$ satisfies (72) if $x_{s}\in\cup_{i=1}^{{L_{D}}}\left({\mathcal{K}(s,T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})\ominus\mathcal{E}_{\text{off-nom}}}\right)$.
See Sec. A.5 for the proof of Prop. 7 using (66) and (67).
We solve (77) using a receding horizon framework. For a finite planning horizon $T_{\text{plan}}\in\mathbb{N}$, the following (non-convex) optimization problem approximates (77),
@@
with $\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})\triangleq\mathcal{K}(k,k+T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})\ominus\mathcal{E}_{\text{off-nom}}$. The non-convexity in (82) arises from the disjunctive constraints [34].
@@
For a sampling time $T_{\text{sample}}=20$ minutes, we solve (77) with a planning horizon of $2$ hours ($T_{\text{plan}}=6$) and an abort-safety time horizon of $6$ hours ($T_{\text{safe}}=18$). We consider the nominal control set $\mathcal{U}=\llbracket-1/3,1/3\rrbracket$ (in m/s), off-nominal input set $\mathcal{U}_{\text{off-nom}}=0.1\mathcal{U}$, post-failure process noise in the ellipsoid $\mathcal{W}_{\text{off-nom}}=(1/60I_{3},0_{3\times 1})$ (in m/s), and post-failure measurement noise in the ellipsoid $\mathcal{E}_{\text{off-nom}}=(\operatorname{diag}(1,1,1,1/60,1/60,1/60),0_{6\times 1})$ (in m and m/s). We define the origin-centered keep-out set $\mathcal{D}={\llbracket-100,100\rrbracket}^{3}$ (in m) which contains the Lunar gateway [19]. We also define a proper cone characterized by four rays originating from the origin as the line-of-sight cone =$\mathcal{L}=\{x\in\mathbb{R}^{3}\mid[0,1,-1;0,1,1;-1,1,0;1,1,0]x\leq 0_{4\times 1}\}$. We define an ellipsoidal target set $\mathcal{T}=(0.05I_{3},c_{\text{target}})$ with $c_{\text{target}}=[0.2416;-0.4017;-0.1738]$ and initial state $x_{0}=[1.4498;-2.4105;-1.0429;0.01;0.01;0.01]$ such that, from the origin, the target is $0.5$ km away and the initial state is $3$ km away with non-zero initial velocity. We rotate $\mathcal{D}$ and $\mathcal{L}$ to have the $+\mathrm{y}$-face of $\mathcal{D}$ and the axis of symmetry of $\mathcal{L}$ be aligned with the line segment joining $x_{0}$ and $c_{\text{target}}$ respectively. The choice of parameters considers a rendezvous approach with the Sun behind the spacecraft, as the gateway flies near the apolune of the NRHO.
@@
The exact computation of $\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})$ based on polytopes [24] is challenging, due to the complexity of the calculations involved in the considered problem setting. Therefore, we use the proposed approach for inner-approximating $T_{\text{safe}}$-RC set using constrained zonotopes to enforce abort-safety constraint. Specifically, we use Thm. 1 and Sec. 5.2 to compute constrained zonotopic, inner-approximations of $\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})$, and then use big-M formulations to cast the disjunctive constraints in (82) as mixed-integer linear constraints.
@@
Consider ${L_{D}}$ constrained zonotopes $\{\mathcal{C}_{i}\}_{i=1}^{{L_{D}}}$ where $\mathcal{C}_{i}\subseteq\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})$ for each $i\in\mathbb{N}_{[1:{L_{D}}]}$ and some $k\in\mathbb{N}_{[t:t+T_{\text{plan}}]}$. Using ${L_{D}}$ auxiliary continuous variables $\xi_{i}\in\mathbb{R}^{{N_{C,i}}}$ and binary variables $\delta_{i}\in\{0,1\}$, and a sufficiently large $M>0$, the following set of $12(n+{N_{C,i}})+{M_{C,i}}+$ 1 mixed-integer linear constraints is sufficient for $x_{k|t}\in\cup_{i=1}^{{L_{D}}}\mathcal{K}_{i}^{\prime}(k,k+T_{\text{safe}})$ at any $k\in\mathbb{N}_{[t+1:t+T_{\text{plan}}]}$,
$\forall i\in\mathbb{N}_{[1:{L_{D}}]},$ $\quad{\|G_{C,i}\xi_{i}+c_{C,i}-x_{k|t}\|}_{\infty}\leq M(1-\delta_{i})$
$\forall i\in\mathbb{N}_{[1:{L_{D}}]},$ $\quad A_{C,i}\xi_{i}=b_{C,i},\ {\|\xi_{i}\|}_{\infty}\leq 1,$
$\quad\sum\nolimits_{i=1}^{{L_{D}}}\delta_{i}\geq 1.$
@@
Similar mixed-integer constraints based on big-M can be used to encode the disjunctive constraint $x_{k|t}\in\cup_{i=1}^{{L_{D}}}\mathcal{H}_{i}^{\complement}$ [34]. Thus, (82) is a mixed-integer quadratic program, which can be solved via off-the-shelf solvers like GUROBI [30].
@@
Figure 6: Designed nominal rendezvous trajectory along with the abort trajectories in case of failures at $t\in\mathbb{N}_{[1:T_{\text{sim}}]}$. The nominal trajectory starts at the initial state (diamond) and reaches the target set $\mathcal{T}$ (green) in $T_{\text{sim}}=10$ time steps, while staying within the line-of-sight cone $\mathcal{L}$ (black). The abort-safety requirement curves the nominal trajectory away from the keep-out set (red) at all times. The abort trajectories stay outside the keep-out set, despite the presence of disturbances which are adversarially chosen according to (85). See https://youtu.be/6BPmHgxD3OI for more details.
@@
We can also compute an abort-safe control $u_{k|t}=\pi_{k}(\hat{z}_{k|t})$ for the current state estimate $\hat{z}_{k|t}$ by solving a convex problem,
$\begin{array}[]{rl}\underset{u_{k|t}\in\mathcal{U}_{\text{off-nom}}}{\text{min.}}&\ J(u_{k|t})\ \text{s.\ t.}\ A_{k}\hat{z}_{k|t}+B_{k}u_{k|t}\in\mathcal{K}_{k+1|t}\ominus\Phi_{k},\end{array}$ (84)
@@
where $J$ is a user-specified, convex cost function on the abort-safe control (we choose $J=\|\cdot\|_{2}$ to minimize post-failure fuel consumption), $\{\mathcal{K}_{k|t}\}_{k=t}^{t+T_{\text{safe}}}$ is the sequence of sets obtained using Sec. 5.2 with $\mathcal{K}_{t+T_{\text{safe}}|t}=\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}}$ and $\mathcal{K}(t,T_{\text{safe}},\mathcal{H}_{i}^{\complement}\ominus\mathcal{E}_{\text{off-nom}})=\mathcal{K}_{t|t}$, as prescribed in Prop. 7. We compute the corresponding adversarial disturbance $\phi_{k|t}$ by solving another convex problem,
$\underset{\phi_{k|t}\in\Phi_{k}}{\text{min}}$ $\quad\operatorname{dist}(A_{k}\hat{z}_{k|t}+B_{k}\pi_{k}(\hat{z}_{k|t})+\phi_{k|t},\mathcal{D}).$ (85)
@@
Here, (85) computes $\phi_{k|t}$ that reduces the distance between the next state estimate and the keep-out set. Thus, we approximate the true worst-case disturbance sequence, whose exact computation would have required solving a two-player game, which is computationally difficult in six dimensions [4].
@@
Fig. 6 demonstrates the designed nominal rendezvous trajectory, which takes $T_{\text{sim}}=10$ time steps ($200$ minutes) to reach $\mathcal{T}$ with a total $\Delta v$ of $1.07$ m/s. The trajectory design required the computation of $385$ $18$-step RC sets, and the set computation took a total of $7.53$ seconds. Fig. 6 also shows the designed abort-safe trajectories originating from each time step of the nominal rendezvous trajectory using the post-failure control (84) and adversarial disturbances (85), along with an outer-approximation of the one-step forward reach set $A_{t}\hat{z}_{k|t}+B_{k|t}\pi_{k}(\hat{z}_{k|t})+\Phi_{k}$. As expected, the abort-safety requirement (72) is satisfied at all times $t\in\mathbb{N}_{[1:T_{\text{sim}}]}$. As the spacecraft approaches the target, we observe that the conservativeness of our enforcement of the abort-safety requirement (72) via Prop. 7 causes the abort trajectories to cluster in front of the keep out set $\mathcal{D}$. In actual rendezvous missions, the spacecraft would receive a go/no-go decision for docking as it nears the target.
@@
Fig. 7 shows the evolution of the distance to the target set (in log-scale) and the cumulative $\Delta v$ expended over the course of the rendezvous, as well as the computation time spent solving (82) at each time step $t\in\mathbb{N}_{[0:T_{\text{sim}}]}$. The rendezvous trajectory initially uses moderately high control inputs $\Delta v$ to steer the spacecraft towards the target while satisfying the abort-safety requirements, and then utilizes the momentum to reach the target set with minimal additional control inputs, as expected. The rendezvous trajectory maintains abort-safety using the constrained zonotope-based constraints computed with the method proposed in this paper.